OpenStudy (anonymous):

Find all zeros of the function and write the polynomial as a product of linear factors. f(x) = x4 + 25x2 + 144 pleasee heeelp!

OpenStudy (anonymous):

do you mean f(x)=4x+194?

OpenStudy (anonymous):

and when the f(x)=0?

OpenStudy (anonymous):

no, thats the problem f(x)= x^4+25x^2+144

OpenStudy (anonymous):

oh, ok i got it sorry my compu's got some probs , i'll try to come back later, hope sb can help caro for now

OpenStudy (anonymous):

Caroo - is there a chance that the middle term is 24x^2 instead of 25x^2?

OpenStudy (harsimran_hs4):

well this seems biquadratic equation but you can rewrite it as y^2 + 25y + 144 = 0 where x^2 = y and solve it just as quadratic equation........i hope you know the formula for this...

OpenStudy (dumbcow):

Factor 16+9 = 25 16*9 = 144 --> (x^2 +16)(x^2 +9) = 0 x^2 +16 = 0 --> x = +-4i x^2 +9 = 0 --> x = +-3i